# Coherence, Transport, and Chaos in 1D Bose-Hubbard Model: Disorder vs. Stark Potential

### v1

@@ Abstract
We study quantum coherence and phase transitions in a finite-size one-dimensional Bose-Hubbard model using exact numerical diagonalization. The system is investigated under the competing effects of thermal fluctuations, a Stark potential, and a disorder term. We compute several observables, including the condensate fraction, superfluid fraction, visibility, number fluctuations, and the $\ell_{1}$-norm of quantum coherence, to characterize the transition from the Mott insulator to the superfluid phase. In the standard Bose-Hubbard model, ground-state properties exhibit signatures of a quantum phase transition under open boundary conditions. While finite-size effects preclude an exact realization of the thermodynamic transition, our findings serve as indicators of the underlying quantum critical behavior of the disordered Bose-Hubbard model. At finite temperatures, this critical point shifts to a smooth crossover, reflecting the suppression of long-range coherence typical of larger systems. A nonzero Stark potential delays this crossover, promoting localization and the formation of non-superfluid condensates. In contrast, thermal fluctuations can induce unexpected coherence through fluctuation-driven tunneling. Disorder disrupts superfluidity but preserves local coherence, with thermal states exhibiting an enhanced $\ell_{1}$-norm of coherence within the Bose glass regime. Our results highlight how disorder, tilt, and temperature jointly reshape the coherence landscape, and offer valuable insights for quantum simulation and the characterization of quantum phases in strongly correlated systems. Non-ergodic behavior is observed in the clean system where the mean gap ratio stays below Poisson and Gaussian orthogonal ensemble (GOE) values. With a Stark potential, level statistics shift from Poisson-like to near-GOE as tunneling increases, which indicates a crossover to semi-ergodic dynamics. In disordered systems, the mean gap ratio starts near Poisson but never reaches GOE. These spectral features support the coherence-based results and reveal how disorder, tilt, and tunneling collectively influence quantum transport.
@@ I Introduction
The Bose-Hubbard model (BHM) has emerged as a cornerstone for understanding quantum phase transitions (QPTs) among cold atoms in optical lattice [1, 2, 3, 4, 5, 2, 6]. At zero temperature, it generally exhibits a transition from superfluid (SF) to Mott insulator (MI), driven by the competition between tunnel energy and site repulsion [4]. This transition, first experimentally observed in ultracold atoms in optical lattices [5], highlights the delicate balance between quantum fluctuations and interactions. However, real-world implementations, whether in ultra-cold gases [2, 7], superconducting circuits [8, 9], or engineered quantum materials [10], must deal with inevitable perturbations: thermal noise, quenched disorder, and Stark potential (SP). Understanding how these factors collectively reshape phase diagram of the BHM is critical for advancing both fundamental many-body physics and the development of robust quantum simulators. In a recent preprint [11], P. M. Clavero and A. Rodríguez identified a chaotic regime in the BHM and mapped the location and extent of this chaotic phase as functions of energy and model parameters. Notably, they found that the chaotic phase of the clean BHM can be enhanced by introducing a moderate tilt. Building on this insight, we investigate the interplay between disorder strength, hopping amplitude, and SP to explore signatures of chaos in the finite-size one-dimensional (1D) disordered BHM.
Disorder [4, 12, 13, 3] and SP [14, 15], both break translational invariance symmetry in the BHM, represent distinct pathways to localization. Quenched disorder, arising from lattice imperfections or inhomogeneous potentials, fragments the system into a Bose glass (BG)—a compressible insulator with rare SF puddles. In contrast, SP, induced by tilting the lattice, suppresses tunneling via Wannier-Stark localization [16], confining bosons to spatially restricted regions. While both mechanisms suppress long-range coherence, their interplay remains poorly understood, as the BHM is generally non-integrable and exhibits chaos in the presence of such potentials.
@@ II The Model
where $\hat{a}_{j}^{\dagger}$ ($\hat{a}_{j}$) creates (annihilates) spinless bosons at site $j$, with $[\hat{a}_{j},\hat{a}_{j^{\prime}}^{\dagger}]=\delta_{jj^{\prime}}$, and $\tau$ denotes the tunneling amplitude. Interactions are introduced via an on-site repulsion term $U$, yielding the clean BHM:
@@
The random site-dependent disorder $\delta_{j}$ is distributed within the range $[-\delta,+\delta]$ with $\delta$ being the energy scale of the strength of the disorder [21]. The total Hamiltonian of the clean BHM, upon inclusion of disorder and the Stark potential $gj$, is extended to:
@@
Figure 1 summarizes the key emergent phases: disorder BG regions [Figs. 1(b) and (c)], while the presence of a SP can localize particles even in the absence of interactions [Figs. 1(d)–(i)]. The thermal state, defined as $\hat{\varrho}_{T}=e^{-\beta\hat{H}}/\mathrm{Tr}(e^{-\beta\hat{H}})$, is obtained from the full energy spectrum $\{|\psi_{i}\rangle,\mathcal{E}_{i}\}$, with each state weighted by a Boltzmann factor $p_{i}\propto e^{-\beta\mathcal{E}_{i}}$. Although simulations are truncated to a system size of $L=6$, essential features such as coherence revivals remain robust across system sizes.
@@ III Quantum Observables and Measurement Metrics
The $f_{c}$, quantifies the fraction of Bose-Einstein condensation according to the Penrose-Onsager criterion [23]. It is defined as the ratio of the largest eigenvalue $\lambda_{\text{max}}$ of the one-body density matrix to the total particle number, namely, $f_{c}={\lambda_{\text{max}}}/{N}$, where the one-body density matrix is defined as $\rho^{(1)}_{ij}=\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle$. For the GS, the one-body correlation function is given by [22] $\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle_{\text{GS}}=\langle\psi_{0}|\hat{a}^{\dagger}_{i}\hat{a}_{j}|\psi_{0}\rangle$, while for TS, it is computed as $\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle_{\text{TS}}=\mathrm{Tr}(\hat{\varrho}_{T}\hat{a}^{\dagger}_{i}\hat{a}_{j})$. In the SF phase, $f_{c}\rightarrow 1$, indicating macroscopic occupation of the lowest momentum state. In contrast, $f_{c}$ can be significantly suppressed in the MI and BG phases. On the other hand, in the SF, $f_{s}$ characterizes the emergence of superfluidity by quantifying the fraction of particles that can flow without viscosity. In the BHM, $f_{s}$ is computed by analyzing the system’s response to a phase twist applied to the hopping terms of the TBH.
@@
The $\mathcal{C}$ provides a basis-dependent quantification of quantum coherence within the density matrix $\rho$, expressed in the Fock basis:
@@ IV.1 Clean BHM
Figure 2 shows key observables versus $\tau/U$ for the clean BHM ($\delta=g=0$), comparing GS (black circles) with TS properties ($T=0.2$, blue squares). The GS energy $\mathcal{E}_{0}$ decreases monotonically with increasing $\tau/U$ as kinetic contributions grow (Fig. 2(a)). The many-body spectrum shown in the inset reveals a gap closing at $\tau/U\approx 0.17$ that hints at the SF-MI QPT.
@@
Interestingly, in the intermediate regime ($0.20.1<\tau/U<$ 0.2), the TS displays slightly enhanced $\mathcal{C}$ and $\mathcal{F}$ compared to GS. This seemingly counterintuitive behavior stems from thermal occupation of low-lying excited states with greater spatial delocalization, which increases local coherence and number fluctuations despite reducing global order.
@@
The level‐spacing analysis of the clean BH chain shows that even as the hopping $\tau/U$ increases from zero, where the system is an MI and the MGR quickly settles to a constant value $\langle r^{\prime}\rangle\approx 0.15$. This value lies well below the Poisson and GOE benchmarks, implying that the spectrum retains non‐ergodic correlations. Such persistent level correlations suggest the presence of approximate conservation laws or hidden symmetries that survive across the MI‐SF crossover.
@@
Figure 4: System-size dependence of the $\langle r^{\prime}\rangle$ versus $\tau/U$ for the clean BHM ($g/U=0.0$, $\mu/U=0.5$, $\delta/U=0.0$) at unit filling. Different system sizes ($L=N=2$ through $L=N=6$) are shown with their respective colored curves. The dashed and dotted horizontal lines represent the theoretical values for GOE (0.53) and Poisson (0.38) statistics, respectively. While the smallest system ($L=N=2$) shows increasing $\langle r^{\prime}\rangle$ with $\tau/U$ that approaches the GOE limit, larger systems maintain consistently low $\langle r^{\prime}\rangle$ values (0.05-0.25) across all hopping strengths, demonstrating persistent non-ergodic spectral correlations that remain stable through the SF-MI transition at $\tau/U\approx 0.17$.
@@ IV.2 Effects of SP
Figure 6 further clarifies the role of the SP by tracking the $\langle r^{\prime}\rangle$ as a function of $\tau/U$. At weak hopping ($\tau/U<0.2$), the system exhibits strong localization with $\langle r^{\prime}\rangle\approx 0.38$, consistent with Poisson statistics. This localized regime arises from the effects of the nonzero SP ($g/U=1.0$), while the chemical potential ($\mu/U=0.5$) modulates the level structure without restoring the delocalization under the unit filling condition. In the crossover regime ($0.60.2<\tau/U<$ 0.6), increasing hopping begins to overcome localization, and $\langle r^{\prime}\rangle$ increases. However, even at large hopping ($\tau/U>0.6$), $\langle r^{\prime}\rangle$ saturates below the GOE value (0.53), reaching only $\approx 0.48$. This indicates partial ergodicity: the system thermalizes only incompletely. These results show that the SP reshapes the ergodic landscape by stabilizing a semi-ergodic regime, where residual localization persists despite strong kinetic energy.
Figure 7 examines the $\langle r^{\prime}\rangle$ versus $\tau/U$ for the BHM with $g=1.0$ across various system sizes at unit filling case. This analysis reveals fundamental changes in spectral statistics induced by the competition between SP and hopping. At small $\tau/U$, all system sizes exhibit $\langle r^{\prime}\rangle$ values near zero, indicating strong spectral correlations consistent with localized states dominated by the SP. As $\tau/U$ increases, we observe a rapid transition toward Poisson statistics ($\langle r^{\prime}\rangle\approx 0.38$) and beyond, with system size significantly influencing this evolution.
@@
The visibility $\mathcal{V}$ (Fig. 8(d)) follows a similar trend, peaking at small $g$ and decreasing at larger values due to loss of long-range interference. The TS visibility remains consistently below the GS due to thermal decoherence.
Besides, the $\ell_{1}$-norm of coherence $\mathcal{C}$ (Fig. 8(e)) exhibits oscillations on a decaying background. These fluctuations result from resonances between neighboring sites at specific $g$ values where inter-site energy differences match tunneling strengths. At these points, coherence temporarily increases despite strong SPs. The TS shows more pronounced coherence peaks, suggesting a thermal population of these resonant states.
Finally, the number fluctuations $\mathcal{F}$ (Fig. 8(f)) reflect the interplay of interactions, tilt, and thermal effects. They decline with increasing SP due to particle localization, while TS maintains larger $\mathcal{F}$ over a broader range, particularly near resonant $g$ values where delocalization reemerges. These results demonstrate that while linear SPs generally induce strong localization and suppress SF and coherence, both GS and TS observables exhibit complex and non-monotonic behaviors with resonance-enhanced coherence even in the strongly localized regime.
@@
To assess the emergence of ergodicity and quantum chaos in the clean BHM, we compute the $\langle r^{\prime}\rangle$ as a function of the $g/U$, as shown in Fig. 10. For the smallest system ($L=N=2$), we observe that $\langle r^{\prime}\rangle$ increases monotonically with $g/U$ and approaches the GOE value $\langle r^{\prime}\rangle\approx 0.53$ at large gradients. This trend suggests the emergence of level repulsion and spectral statistics characteristic of quantum chaotic systems. However, this behavior is not sustained at larger system sizes. For $L=3$, a weak maximum in $\langle r^{\prime}\rangle$ is seen near $g/U\sim 5$, but its value remains below the GOE limit and subsequently decreases. As the system size increases ($L=4,5,6$), the values of $\langle r^{\prime}\rangle$ remain significantly below the GOE threshold and approach or fall below the Poisson value $\langle r^{\prime}\rangle\approx 0.38$. This trend indicates a lack of level repulsion and is consistent with integrable or localized behavior. Despite the breaking of translational symmetry by the linear gradient, the SP does not induce a sufficient level of mixing to generate chaotic spectral statistics at these sizes and parameters. Instead, the system retains a nonergodic character throughout the range of $g/U$ studied.
@@ IV.3 Role of Disorder
The level statistics show markedly different behavior compared to the clean system case. While $\langle r^{\prime}\rangle$ begins near the Poisson value $0.38$ at small $\tau/U$ characteristic of localized states in the strong disorder regime - it fails to reach the GOE limit $0.53$ even at maximal $\tau/U$ values. This incomplete transition toward quantum chaos suggests that the disorder $\delta/U=1.0$ creates an intermediate regime where neither full localization nor complete ergodicity is achieved.
@@
These results demonstrate how disorder ($\delta=1.0$) can suppress the development of quantum chaos even in the presence of significant hopping ($\tau/U\sim 0.8$), suggesting potential pathways for stabilizing non-thermal states in disordered bosonic systems. The intermediate $\langle r^{\prime}\rangle$ values may indicate a Griffiths-like regime where rare regions dominate the system’s behavior.
Figure 13 illustrates the MGR $\langle r^{\prime}\rangle$ versus normalized hopping $\tau/U$ for systems with quenched disorder ($\delta=1.0$) across different system sizes while maintaining unit filling. This analysis provides a revealing contrast to our findings for systems with SP (Fig. 11) and clean systems.
@@
Figure 14: Effect of disorder strength $\delta/U$ on the BHM observables with zero SP ($g/U=0$). GS (black circles) and TS at $T/U=0.1$ (blue squares). (a) GS energy $\mathcal{E}_{0}$ with spectrum inset, (b) $f_{c}$ (c) $f_{s}$, (d) $\mathcal{V}$, (e) $\mathcal{C}$, and (f) $\mathcal{F}$.
@@
All coherence metrics—$f_{c}$ (Fig. 14(b)), $f_{s}$ (Fig. 14(c)), $\mathcal{V}$ (Fig. 14(d)), and $\mathcal{C}$ (Fig. 14(e))—decrease monotonically with increasing $\delta$, confirming disorder-induced localization. The SF fraction shows the most dramatic reduction, dropping to near-zero in moderate disorder ($\delta\gtrsim 1.0$), while $f_{c}$ decreases more slowly. This difference highlights that condensation can persist in disconnected regions even when global superfluidity is destroyed. Notably, the TS consistently shows enhanced coherence $\mathcal{C}$ and number fluctuations $\mathcal{F}$ (Fig. 14(f)) compared to the GS across all disorder strengths. This thermal enhancement becomes proportionally larger at stronger disorder, where thermal excitations activate tunneling between otherwise disconnected localized states, creating short-range coherence absent in the GS.
@@
To characterize the onset of localization, we analyze the spectral statistics of the disordered BH Hamiltonian by computing the full many-body energy spectrum and by evaluating the adjacent level spacing ratio $\langle r^{\prime}\rangle$, a standard probe of ergodicity breaking. As shown in Fig. 15, the eigenenergy spectrum (top panel) exhibits increased irregularity and bandwidth with increasing disorder strength $\delta/U$, consistent with the emergence of localization. The average ratio $\langle r^{\prime}\rangle$ (bottom panel) transitions from values near the GOE prediction ($\langle r^{\prime}\rangle\approx 0.53$), indicative of ergodic level repulsion, toward the Poissonian limit ($\langle r^{\prime}\rangle\approx 0.38$), characteristic of localized, nonthermal spectra. This crossover, occurring around $\delta/U\approx 1$, signals a breakdown of spectral correlations and provides strong evidence for many-body localization in this interacting system at finite size.
@@
Figure 16 illustrates the MGR $\langle r^{\prime}\rangle$ versus normalized hopping $\tau/U$ for systems with quenched disorder ($\delta/U=1.0$) across different system sizes while maintaining unit filling. This analysis contrasts with our findings for systems with SP (Figure 11) and clean systems. The smallest system with $L=N=2$ (blue curve) exhibits striking behavior, with $\langle r^{\prime}\rangle$ starting near Poisson statistics ($\approx 0.07$) at small $\tau/U$ before sharply rising to a pronounced peak of approximately 0.98 near $\tau/U\approx 2.2$. The MGR then gradually decreases and stabilizes around 0.85 at large hopping strengths. This extreme super-GOE behavior indicates a strong repulsion in this minimal system whens disorder and significant hopping are present.

### v2

@@ Abstract
We study quantum coherence and phase transitions in a finite one-dimensional Bose-Hubbard model using exact diagonalization under thermal fluctuations, a Stark potential, and disorder. We compute the condensate fraction, superfluid fraction, visibility, number fluctuations, and the $\ell_{1}$-norm of coherence to characterize the Mott insulator–superfluid transition. While finite-size effects prevent a sharp transition, ground-state properties reveal signatures of quantum criticality. Thermal fluctuations can enhance coherence via tunneling, a Stark potential promotes localization, and disorder suppresses global superfluidity while preserving local coherence. Our results highlight how disorder, tilt, and temperature reshape coherence and offer insights for quantum simulation and strongly correlated phases. For systems up to six sites with unit filling, we also perform a spectral analysis through the metric mean gap ratio (MGR). However, limited statistics due to the small system size and computational constraints prevent a complete characterization of quantum chaos, yielding only approximate signatures.
@@ I Introduction
The Bose-Hubbard model (BHM) has emerged as a cornerstone for understanding quantum phase transitions (QPTs) among cold atoms in optical lattice [1, 2, 3, 4, 5, 2, 6]. At zero temperature, it generally exhibits a transition from superfluid (SF) to Mott insulator (MI), driven by the competition between tunnel energy and site repulsion [4]. This transition, first experimentally observed in ultracold atoms in optical lattices [5], highlights the delicate balance between quantum fluctuations and interactions. However, real-world implementations, whether in ultra-cold gases [2, 7], superconducting circuits [8, 9], or engineered quantum materials [10], must deal with inevitable perturbations: thermal noise, quenched disorder, and Stark potential (SP). Understanding how these factors collectively reshape phase diagram of the BHM is critical for advancing both fundamental many-body physics and the development of robust quantum simulators. In a recent preprint [11], P. M. Clavero and A. Rodríguez identified a chaotic regime in the BHM and mapped the location and extent of this chaotic phase as functions of energy and model parameters. Notably, they found that the chaotic phase of the clean BHM can be enhanced by introducing a moderate tilt. Building on this insight, we investigate the interplay between disorder strength, hopping amplitude, and SP to explore signatures of chaos in the finite-size one-dimensional (1D) disordered BHM [12, 13].
Disorder [4, 14, 15, 3, 16] and SP [17, 18], both break translational invariance symmetry in the BHM, represent distinct pathways to localization. Quenched disorder, arising from lattice imperfections or inhomogeneous potentials, fragments the system into a Bose glass (BG)—a compressible insulator with rare SF puddles. In contrast, SP, induced by tilting the lattice, suppresses tunneling via Wannier-Stark localization [19], confining bosons to spatially restricted regions. While both mechanisms suppress long-range coherence, their interplay remains poorly understood, as the BHM is generally non-integrable and exhibits chaos in the presence of such potentials.
@@ II The Model
where $\hat{a}_{j}^{\dagger}$ ($\hat{a}_{j}$) creates (annihilates) spinless bosons at site $j$, with $\hat{a}_{j},\hat{a}_{j^{\prime}}^{\dagger}]=\delta_{jj^{\prime}}$, and $\tau$ denotes the tunneling amplitude. Interactions are introduced via an on-site repulsion term $U$, yielding the clean BHM:
@@
The random site-dependent disorder $\delta_{j}$ is distributed within the range $\delta,+\delta]$ with $\delta$ being the energy scale of the strength of the disorder [24]. The total Hamiltonian of the clean BHM, upon inclusion of disorder and the Stark potential $gj$, is extended to:
@@
[…]
@@ III Quantum Observables and Measurement Metrics
The condensate fraction $f_{c}$ quantifies the fraction of Bose-Einstein condensation according to the Penrose-Onsager criterion [26]. It is defined as the ratio of the largest eigenvalue $\lambda_{\text{max}}$ of the one-body density matrix to the total particle number, namely, $f_{c}={\lambda_{\text{max}}}/{N}$, where the one-body density matrix is defined as $\rho^{(1)}_{ij}=\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle$. For the GS, the one-body correlation function is given by [25] $\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle_{\text{GS}}=\langle\psi_{0}|\hat{a}^{\dagger}_{i}\hat{a}_{j}|\psi_{0}\rangle$, while for TS, it is computed as $\langle\hat{a}^{\dagger}_{i}\hat{a}_{j}\rangle_{\text{TS}}=\mathrm{Tr}(\hat{\varrho}_{T}\hat{a}^{\dagger}_{i}\hat{a}_{j})$. In the SF phase, $f_{c}\rightarrow 1$, indicating macroscopic occupation of the lowest momentum state. In contrast, $f_{c}$ can be significantly suppressed in the MI and BG phases. On the other hand, in the SF, $f_{s}$ characterizes the emergence of superfluidity by quantifying the fraction of particles that can flow without viscosity. In the BHM, $f_{s}$ is computed by analyzing the system’s response to a phase twist applied to the hopping terms of the TBH.
@@
The following formula provides a basis-dependent quantification of quantum coherence within the density matrix $\rho$, expressed in the Fock basis:
@@ IV.1 Clean BHM
Figure 2 shows key observables versus $\tau/U$ for the clean BHM ($\delta/U=g/U=0$), comparing GS (black circles) with TS properties ($T/U=0.2$, blue squares). The GS energy $\mathcal{E}_{0}$ decreases monotonically with increasing $\tau/U$ as kinetic contributions grow (Fig. 2(a)). The many-body spectrum shown in the inset reveals a gap closing at $\tau/U\approx 0.17$ that hints at the SF-MI QPT.
@@
Interestingly, in the intermediate regime ($0.1<\tau/U<0.2$), the TS displays slightly enhanced $\mathcal{C}$ and $\mathcal{F}$ compared to GS. This seemingly counterintuitive behavior stems from thermal occupation of low-lying excited states with greater spatial delocalization, which increases local coherence and number fluctuations despite reducing global order.
@@
The level‐spacing analysis of the clean BH chain shows that even as the hopping $\tau/U$ increases from zero, where the system is an MI, the MGR quickly settles to a constant value $\langle r^{\prime}\rangle\approx 0.15$ (see Fig. 3). This value lies well below the Poisson and GOE benchmarks, implying that the spectrum retains non‐ergodic correlations. Such persistent level correlations suggest the presence of approximate conservation laws or hidden symmetries that survive across the MI‐SF crossover.
@@
Figure 4: System-size dependence of the $\langle r^{\prime}\rangle$ versus $\tau/U$ for the clean BHM ($g/U=0.0$, $\mu/U=0.5$, $\delta/U=0.0$) at unit filling. Different system sizes ($L=N=2$ through $L=N=6$) are shown with their respective colored curves. The dashed and dotted horizontal lines represent the theoretical values for GOE (0.53) and Poisson (0.38) statistics, respectively. While the smallest system ($L=N=2$) shows increasing $\langle r^{\prime}\rangle$ with $\tau/U$ that approaches the GOE limit, larger systems maintain consistently low $\langle r^{\prime}\rangle$ values (0.05-0.25) across all hopping strengths, demonstrating persistent non-ergodic spectral correlations that remain stable through the SF-MI transition at $\tau/U\approx 0.17$. For the minimal two-site case ($L=N=2$), the apparent GOE-like value of the mean gap ratio $\langle r^{\prime}\rangle$ is a finite-size artifact rather than genuine chaos. With only three many-body states, the model is integrable, but poor spectral statistics, boundary effects, and kinetic-energy dominance mimic level repulsion. For larger systems ($L\geq 3$), these artifacts disappear and $\langle r^{\prime}\rangle$ settles at sub-Poissonian values, consistent with nonergodic behavior.
@@ IV.2 Effects of SP
Figure 6 further clarifies the role of the SP by tracking the $\langle r^{\prime}\rangle$ as a function of $\tau/U$. At weak hopping ($\tau/U<0.2$), the system exhibits strong localization with $\langle r^{\prime}\rangle\approx 0.38$, consistent with Poisson statistics. This localized regime arises from the effects of the nonzero SP ($g/U=1.0$), while the chemical potential ($\mu/U=0.5$) modulates the level structure without restoring the delocalization under the unit filling condition. In the crossover regime ($0.2<\tau/U<0.6$), increasing hopping begins to overcome localization, and $\langle r^{\prime}\rangle$ increases. However, even at large hopping ($\tau/U>0.6$), $\langle r^{\prime}\rangle$ saturates below the GOE value (0.53), reaching only $\approx 0.48$. This indicates partial ergodicity: the system thermalizes only incompletely. These results show that the SP reshapes the ergodic landscape by stabilizing a semi-ergodic regime, where residual localization persists despite strong kinetic energy.
Figure 7 examines the $\langle r^{\prime}\rangle$ versus $\tau/U$ for the BHM with $g/U=1.0$ across various system sizes at unit filling case. This analysis reveals fundamental changes in spectral statistics induced by the competition between SP and hopping. At small $\tau/U$, all system sizes exhibit $\langle r^{\prime}\rangle$ values near zero, indicating strong spectral correlations consistent with localized states dominated by the SP. As $\tau/U$ increases, we observe a rapid transition toward Poisson statistics ($\langle r^{\prime}\rangle\approx 0.38$) and beyond, with system size significantly influencing this evolution.
@@
The visibility $\mathcal{V}$ (Fig. 8(d)) follows a similar trend, peaking at small $g/U$ and decreasing at larger values due to loss of long-range interference. The TS visibility remains consistently below the GS due to thermal decoherence.
Besides, the $\ell_{1}$-norm of coherence $\mathcal{C}$ (Fig. 8(e)) exhibits oscillations on a decaying background. These fluctuations result from resonances between neighboring sites at specific $g/U$ values where inter-site energy differences match tunneling strengths. At these points, coherence temporarily increases despite strong SPs. The TS shows more pronounced coherence peaks, suggesting a thermal population of these resonant states.
Finally, the number fluctuations $\mathcal{F}$ (Fig. 8(f)) reflect the interplay of interactions, tilt, and thermal effects. They decline with increasing SP due to particle localization, while TS maintains larger $\mathcal{F}$ over a broader range, particularly near resonant $g/U$ values where delocalization reemerges. These results demonstrate that while linear SPs generally induce strong localization and suppress SF and coherence, both GS and TS observables exhibit complex and non-monotonic behaviors with resonance-enhanced coherence even in the strongly localized regime.
@@
To assess the emergence of ergodicity and quantum chaos in the clean BHM, we compute the $\langle r^{\prime}\rangle$ as a function of the $g/U$, as shown in Fig. 10. For the smallest system ($L=N=2$), we observe that $\langle r^{\prime}\rangle$ increases with $g/U$ and approaches the GOE value $\langle r^{\prime}\rangle\approx 0.53$ at large gradients. This trend suggests the emergence of level repulsion and spectral statistics characteristic of quantum chaotic systems. However, this behavior is not sustained at larger system sizes. For $L=3$, a weak maximum in $\langle r^{\prime}\rangle$ is seen near $g/U\sim 5$, but its value remains below the GOE limit and subsequently decreases. As the system size increases ($L=4,5,6$), the values of $\langle r^{\prime}\rangle$ remain significantly below the GOE threshold and approach or fall below the Poisson value $\langle r^{\prime}\rangle\approx 0.38$. This trend indicates a lack of level repulsion and is consistent with integrable or localized behavior. Despite the breaking of translational symmetry by the linear gradient, the SP does not induce a sufficient level of mixing to generate chaotic spectral statistics at these sizes and parameters. Instead, the system retains a nonergodic character throughout the range of $g/U$ studied.
@@ IV.3 Role of Disorder
The level statistics show markedly different behavior compared to the clean system case. $\langle r^{\prime}\rangle$ begins near the Poisson value $0.38$ in small $\tau/U$, characteristic of localized states in the strong disorder regime, and reaches the GOE limit $0.53$ at some values of $\tau/U$. This incomplete transition toward quantum chaos suggests that the disorder $\delta/U=1.0$ creates an intermediate regime where neither full localization nor complete ergodicity is achieved.
@@
These results demonstrate how disorder ($\delta/U=1.0$) can suppress the development of quantum chaos even in the presence of significant hopping ($\tau/U\sim 0.8$), suggesting potential pathways for stabilizing non-thermal states in disordered bosonic systems. The intermediate $\langle r^{\prime}\rangle$ values may indicate a Griffiths-like regime where rare regions dominate the system’s behavior.
Figure 13 illustrates the MGR $\langle r^{\prime}\rangle$ versus normalized hopping $\tau/U$ for systems with quenched disorder ($\delta/U=1.0$) across different system sizes while maintaining unit filling. This analysis provides a revealing contrast to our findings for systems with SP (Fig. 11) and clean systems.
@@
Figure 14: Effect of disorder strength $\delta/U$ on the BHM observables with zero SP ($g/U=0$). GS (black circles) and TS at $T/U=0.1$ (blue squares). (a) GS energy $\mathcal{E}_{0}$ with spectrum inset, (b) $f_{c}$, (c) $f_{s}$, (d) $\mathcal{V}$, (e) $\mathcal{C}$, and (f) $\mathcal{F}$.
@@
[…]
@@
To characterize the onset of localization, we analyze the spectral statistics of the disordered BH Hamiltonian by computing the full many-body energy spectrum and evaluating the adjacent level spacing ratio $\langle r^{\prime}\rangle$, a standard probe of ergodicity breaking. As shown in Fig. 15, the eigenenergy spectrum (top panel) exhibits increased irregularity and bandwidth with increasing disorder strength $\delta/U$, consistent with the emergence of localization. The average ratio $\langle r^{\prime}\rangle$ (bottom panel) transitions from values near the GOE prediction ($\langle r^{\prime}\rangle\approx 0.53$), indicative of ergodic level repulsion, toward the Poissonian limit ($\langle r^{\prime}\rangle\approx 0.38$), characteristic of localized, nonthermal spectra. This crossover, occurring around $\delta/U\approx 1$, signals a breakdown of spectral correlations and provides strong evidence for many-body localization in this interacting system at finite size.
@@
Figure 16 illustrates MGR $\langle r^{\prime}\rangle$ versus normalized hopping $\tau/U$ for systems with quenched disorder ($\delta/U=1.0$) across different system sizes while maintaining unit filling. This analysis contrasts with our findings for systems with SP (Figure 11) and clean systems. The smallest system with $L=N=2$ (blue curve) exhibits striking behavior, with $\langle r^{\prime}\rangle$ starting near Poisson statistics ($\approx 0.07$) at small $\tau/U$ before sharply rising to a pronounced peak of approximately 0.98 near $\tau/U\approx 2.2$. The MGR then gradually decreases and stabilizes around 0.85 at large hopping strengths. This extreme super-GOE behavior indicates a strong repulsion in this minimal system when disorder and significant hopping are present.